1000 gm aqueous solution of caco3 contains 10g of calcium carbonate concentration of the solution

Grade:12th pass

1 Answers

Himanshu Rathour
5 years ago
With the given quantities we can find only 2 types of conc
1)w/w
w/w conc=10/1000=0.01
2)Molality
m=moles ofoCaCO3/wt of solvent
=0.1/1
m=0.1